# Lipschitz-free spaces and Bossard’s reduction argument

### v2

@@ 3. Three main results
The map $\Xi$ is continuous. Indeed, let $d\in\mathcal{R}$ and $v_{1},\ldots,v_{k}\in V$. There exists $N\in\omega$ such that $v_{\ell}(i)=0$ whenever $\ell=1,\ldots,k$ and $i\geqslant N$. Consider any $d^{\prime}\in\mathcal{R}$ such that $d^{\prime}(i,j)=d(i,j)$ whenever $i,j\leqslant N$. Then $\Xi(d^{\prime})(v_{\ell})=\Xi(d)(v_{\ell})$ for $\ell=1,\ldots,k$. This holds because, given $f\in B_{\operatorname{Lip}_{0}(M_{d^{\prime}})}$, by McShane’s extension theorem [weaver:18]*Theorem 1.33 there exists $g\in B_{\operatorname{Lip}_{0}(M_{d})}$ such that $g(i)=f(i)$ whenever $i\leqslant N$, and vice-versa. Therefore $\Xi$ is continuous. ∎
@@
Evidently $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}=\Xi^{-1}(\langle{X}\rangle^{\mathcal{B}}_{\hookrightarrow})$ is $\mathbf{\Sigma}^{1}_{1}$. Similarly to [clemens:12]*Lemma 4, the set
@@
Using the Borel isomorphism ${\Theta}:{\mathcal{B}}\to{\mathcal{SB}}$ and [bossard:02]*Corollary 3.3, the set of $\mu\in\mathcal{B}$ such that $X_{\mu}$ has the RNP is $\mathbf{\Pi}^{1}_{1}$. Now apply Proposition 3.1 and [aliaga:gartand:petitjean:prochazka:22]*Theorem 4.6. ∎
@@
According to [kechris:95]*Exercise 23.1, $P$ is $\mathbf{\Pi}^{0}_{3}$-complete. Now set $\Gamma=\omega\times\omega$, $M=\ell_{1}(\Gamma\cup\{\infty\})$, $A=\{0\}$ and $\phi_{0}=\{(0,0)\}$. Let $d$ be the usual metric on $M$, let ${\tau}:{\omega\setminus A}\to{\Gamma}$ be a bijection, let $(e_{\gamma})_{\gamma\in\Gamma\cup\{\infty\}}$ be the usual basis, and define injections ${\phi_{1},\phi_{2}}:{\omega\setminus A}\to{M}$ by
@@
The final lemma below uses Bossard’s reduction argument [bossard:02]*Theorem 3.2; Theorems 1.11.3 emerge as relatively straightforward consequences of it.
@@
Let $X$ be a separable Banach space. By Proposition 3.2, $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. Let $d^{\prime}\in\mathcal{R}$ and $d\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$ such that $M_{d^{\prime}}$ is isometric to a 1-Lipschitz retract of $M_{d}$. By [weaver:18]*Theorems 3.6 and 3.7, $\mathcal{F}({M_{d^{\prime}}})\hookrightarrow\mathcal{F}({M_{d}})$ and thus $d^{\prime}\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$.
Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$. By Lemma 3.10, $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}=\mathcal{R}$. As $Y\hookrightarrow\mathcal{F}({c_{0}})$ whenever $Y$ is a separable Banach space ([dutrieux:ferenczi:06]*p. 1042) and $c_{0}\equiv M_{d}$ for some $d\in\mathcal{R}$, it follows that $Y\hookrightarrow X$ for all such $Y$. ∎
@@
Let $M$ be a complete separable metric space. By Proposition 3.2, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$. By Lemma 3.10, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}=\mathcal{R}$. By considering e.g. $\mathbb{U}$ or $c_{0}$ (the latter being Lipschitz universal by Aharoni’s theorem, see e.g. [benyamini:lindenstrauss:00]*Theorem 7.11), similarly to above we see that $M$ is also Lipschitz universal. ∎
@@
Using [ghawadrah:17], [cuth:doucha:kurka:22a]*Theorem 8 and Proposition 3.1, the set
@@
is $\mathbf{\Delta}^{1}_{1}$. Let $d^{\prime}\in\mathcal{R}$ and $d\in\mathcal{R}_{\mathrm{BAP}}$ such that $M_{d^{\prime}}$ is isometric to a 1-Lipschitz retract of $M_{d}$. Again by [weaver:18]*Theorems 3.6 and 3.7, $\mathcal{F}({M_{d^{\prime}}})$ is linearly isometric to a 1-complemented subspace of $\mathcal{F}({M_{d}})$ and thus $d^{\prime}\in\mathcal{R}_{\mathrm{BAP}}$. For a contradiction assume $\mathcal{R}_{\mathrm{dis}}\subseteq\mathcal{R}_{\mathrm{BAP}}$. Then $\mathcal{R}_{\mathrm{BAP}}=\mathcal{R}$ by Lemma 3.10. However, this is false because we can pick $d\in\mathcal{R}$ such that $M_{d}$ is isometric to a separable Banach space that fails the AP; as is pointed out in the Introduction, $d\notin\mathcal{R}_{\mathrm{BAP}}$ by [godefroy:kalton:03]*Theorem 3.1.
Now pick $d\in\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$. Again, as stated in the Introduction, $\mathcal{F}({M_{d}})$ has the AP by [aliaga:nous:petitjean:prochazka:21]*Corollary 2.8. That $\mathcal{F}({M_{d}})$ cannot be isomorphic to a dual space is a simple consequence of the classical result of Grothendieck which asserts that a separable Banach space has the MAP whenever it has the AP and is isometric to a dual space. ∎
@@
    Given metric spaces $M$ and $N$, let us write $M\stackrel{{\scriptstyle u}}{{\hookrightarrow}}N$ if $M$ is uniformly homeomorphic to a subspace of $N$. Similarly to [clemens:12]*Lemma 4, the set
@@
    We can replace $\mathcal{R}_{\mathrm{dis}}$ by $\mathcal{R}_{\mathrm{p1u}}$ to show that the set $\mathcal{R}_{\mathrm{p1u}}\setminus\mathcal{R}_{\mathrm{BAP}}$ (which is $\mathbf{\Pi}^{1}_{1}$ by Proposition 3.4) is not $\mathbf{\Delta}^{1}_{1}$. No changes to the proofs above are needed. Given $d\in\mathcal{R}_{\mathrm{p1u}}$, we have $\langle\mathcal{F}({M_{d}})\rangle^{\mathcal{R}}_{\hookrightarrow}\subseteq\mathcal{R}_{\mathrm{p1u}}$ by [aliaga:gartand:petitjean:prochazka:22]*Theorem 4.6. Therefore, by arguing as above, there is a sequence $(d_{\alpha})_{\alpha<\omega_{1}}$ of elements of $\mathcal{R}_{\mathrm{p1u}}\setminus\mathcal{R}_{\mathrm{BAP}}$ such that $\mathcal{F}({M_{d_{\beta}}})\not\hookrightarrow\mathcal{F}({M_{d_{\alpha}}})$ whenever $\alpha<\beta<\omega_{1}$. I don’t know if the spaces $\mathcal{F}({M_{d_{\alpha}}})$ have the AP.
@@
Apply Lemma 3.9 to a separable Banach space $M$ that fails the AP. If $T\in\mathbf{WF}$ then $\mathcal{F}({M_{\Delta(T)}})$ has the AP again by [aliaga:nous:petitjean:prochazka:21]*Corollary 2.8. If $T\in\mathbf{IF}$ then $M$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$. Once again by [weaver:18]*Theorems 3.6 and 3.7, $\mathcal{F}({M})$ is linearly isometric to a 1-complemented subspace of $\mathcal{F}({M_{\Delta(T)}})$. Since $\mathcal{F}({M})$ fails the AP, so must $\mathcal{F}({M_{\Delta(T)}})$. Therefore $\Delta$ witnesses the fact that $\mathbf{WF}$ is reducible to $\mathcal{R}_{\mathrm{AP}}$. ∎
@@ 4. Complexity and the approximation property
This section is solely concerned with determining the complexity of two classes of Banach spaces having the AP. The proof of the result below is an adaptation of an argument from [ghawadrah:17]. As above, we frame the result in terms of the Polish space $\mathcal{B}$ instead of the Effros-Borel space $\mathcal{SB}$.
@@
Hence we can define a bounded $\mathbb{Q}$-linear operator ${T}:{V}\to{\operatorname{span}(e_{i})_{i<R}}$ by setting $Te_{i}=w_{i}$, $i\in\omega$, and then extending to finite $\mathbb{Q}$-linear combinations of the $e_{i}$. By equicontinuity we extend it again to a bounded linear operator on $X_{\mu}$ satisfying $TX_{\mu}\subseteq\operatorname{span}(e_{i})_{i<R}$ and $\mu(T)\leqslant L$.

### v3

@@ 3. Three main results
The map $\Xi$ is continuous. Indeed, let $d\in\mathcal{R}$ and $v_{1},\ldots,v_{k}\in V$. There exists $N\in\omega$ such that $v_{\ell}(i)=0$ whenever $\ell=1,\ldots,k$ and $i\geqslant N$. Consider any $d^{\prime}\in\mathcal{R}$ such that $d^{\prime}(i,j)=d(i,j)$ whenever $i,j\leqslant N$. Then $\Xi(d^{\prime})(v_{\ell})=\Xi(d)(v_{\ell})$ for $\ell=1,\ldots,k$. This holds because, given $f\in B_{\operatorname{Lip}_{0}(M_{d^{\prime}})}$, by McShane’s extension theorem [31]*Theorem 1.33 there exists $g\in B_{\operatorname{Lip}_{0}(M_{d})}$ such that $g(i)=f(i)$ whenever $i\leqslant N$, and vice-versa. Therefore $\Xi$ is continuous. ∎
@@
Evidently $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}=\Xi^{-1}(\langle{X}\rangle^{\mathcal{B}}_{\hookrightarrow})$ is $\mathbf{\Sigma}^{1}_{1}$. Similarly to [8]*Lemma 4, the set
@@
Using the Borel isomorphism ${\Theta}:{\mathcal{B}}\to{\mathcal{SB}}$ and [6]*Corollary 3.3, the set of $\mu\in\mathcal{B}$ such that $X_{\mu}$ has the RNP is $\mathbf{\Pi}^{1}_{1}$. Now apply Proposition 3.1 and [1]*Theorem 4.6. ∎
@@
According to [23]*Exercise 23.1, $P$ is $\mathbf{\Pi}^{0}_{3}$-complete. Now set $\Gamma=\omega\times\omega$, $M=\ell_{1}(\Gamma\cup\{\infty\})$, $A=\{0\}$ and $\phi_{0}=\{(0,0)\}$. Let $d$ be the usual metric on $M$, let ${\tau}:{\omega\setminus A}\to{\Gamma}$ be a bijection, let $(e_{\gamma})_{\gamma\in\Gamma\cup\{\infty\}}$ be the usual basis, and define injections ${\phi_{1},\phi_{2}}:{\omega\setminus A}\to{M}$ by
@@
The final lemma below uses Bossard’s reduction argument [6]*Theorem 3.2; Theorems 1.11.3 emerge as relatively straightforward consequences of it.
@@
Let $X$ be a separable Banach space. By Proposition 3.2, $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. Let $d^{\prime}\in\mathcal{R}$ and $d\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$ such that $M_{d^{\prime}}$ is isometric to a 1-Lipschitz retract of $M_{d}$. By [31]*Theorems 3.6 and 3.7, $\mathcal{F}({M_{d^{\prime}}})\hookrightarrow\mathcal{F}({M_{d}})$ and thus $d^{\prime}\in\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$.
Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$. By Lemma 3.10, $\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}=\mathcal{R}$. As $Y\hookrightarrow\mathcal{F}({c_{0}})$ whenever $Y$ is a separable Banach space ([12]*p. 1042) and $c_{0}\equiv M_{d}$ for some $d\in\mathcal{R}$, it follows that $Y\hookrightarrow X$ for all such $Y$. ∎
@@
Let $M$ be a complete separable metric space. By Proposition 3.2, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}$ is $\mathbf{\Sigma}^{1}_{1}$. Assume that $\mathcal{R}_{\mathrm{dis}}\subseteq\langle X\rangle^{\mathcal{R}}_{\hookrightarrow}$. By Lemma 3.10, $\langle M\rangle^{\mathcal{R}}_{\hookrightarrow}=\mathcal{R}$. By considering e.g. $\mathbb{U}$ or $c_{0}$ (the latter being Lipschitz universal by Aharoni’s theorem, see e.g. [5]*Theorem 7.11), similarly to above we see that $M$ is also Lipschitz universal. ∎
@@
Using [17], [10]*Theorem 8 and Proposition 3.1, the set
@@
is $\mathbf{\Delta}^{1}_{1}$. Let $d^{\prime}\in\mathcal{R}$ and $d\in\mathcal{R}_{\mathrm{BAP}}$ such that $M_{d^{\prime}}$ is isometric to a 1-Lipschitz retract of $M_{d}$. Again by [31]*Theorems 3.6 and 3.7, $\mathcal{F}({M_{d^{\prime}}})$ is linearly isometric to a 1-complemented subspace of $\mathcal{F}({M_{d}})$ and thus $d^{\prime}\in\mathcal{R}_{\mathrm{BAP}}$. For a contradiction assume $\mathcal{R}_{\mathrm{dis}}\subseteq\mathcal{R}_{\mathrm{BAP}}$. Then $\mathcal{R}_{\mathrm{BAP}}=\mathcal{R}$ by Lemma 3.10. However, this is false because we can pick $d\in\mathcal{R}$ such that $M_{d}$ is isometric to a separable Banach space that fails the AP; as is pointed out in the Introduction, $d\notin\mathcal{R}_{\mathrm{BAP}}$ by [18]*Theorem 3.1.
Now pick $d\in\mathcal{R}_{\mathrm{dis}}\setminus\mathcal{R}_{\mathrm{BAP}}$. Again, as stated in the Introduction, $\mathcal{F}({M_{d}})$ has the AP by [2]*Corollary 2.8. That $\mathcal{F}({M_{d}})$ cannot be isomorphic to a dual space is a simple consequence of the classical result of Grothendieck which asserts that a separable Banach space has the MAP whenever it has the AP and is isometric to a dual space. ∎
@@
    Given metric spaces $M$ and $N$, let us write $M\stackrel{{\scriptstyle u}}{{\hookrightarrow}}N$ if $M$ is uniformly homeomorphic to a subspace of $N$. Similarly to [8]*Lemma 4, the set
@@
    We can replace $\mathcal{R}_{\mathrm{dis}}$ by $\mathcal{R}_{\mathrm{p1u}}$ to show that the set $\mathcal{R}_{\mathrm{p1u}}\setminus\mathcal{R}_{\mathrm{BAP}}$ (which is $\mathbf{\Pi}^{1}_{1}$ by Proposition 3.4) is not $\mathbf{\Delta}^{1}_{1}$. No changes to the proofs above are needed. Given $d\in\mathcal{R}_{\mathrm{p1u}}$, we have $\langle\mathcal{F}({M_{d}})\rangle^{\mathcal{R}}_{\hookrightarrow}\subseteq\mathcal{R}_{\mathrm{p1u}}$ by [1]*Theorem 4.6. Therefore, by arguing as above, there is a sequence $(d_{\alpha})_{\alpha<\omega_{1}}$ of elements of $\mathcal{R}_{\mathrm{p1u}}\setminus\mathcal{R}_{\mathrm{BAP}}$ such that $\mathcal{F}({M_{d_{\beta}}})\not\hookrightarrow\mathcal{F}({M_{d_{\alpha}}})$ whenever $\alpha<\beta<\omega_{1}$. I don’t know if the spaces $\mathcal{F}({M_{d_{\alpha}}})$ have the AP.
@@
Apply Lemma 3.9 to a separable Banach space $M$ that fails the AP. If $T\in\mathbf{WF}$ then $\mathcal{F}({M_{\Delta(T)}})$ has the AP again by [2]*Corollary 2.8. If $T\in\mathbf{IF}$ then $M$ is isometric to a 1-Lipschitz retract of $M_{\Delta(T)}$. Once again by [31]*Theorems 3.6 and 3.7, $\mathcal{F}({M})$ is linearly isometric to a 1-complemented subspace of $\mathcal{F}({M_{\Delta(T)}})$. Since $\mathcal{F}({M})$ fails the AP, so must $\mathcal{F}({M_{\Delta(T)}})$. Therefore $\Delta$ witnesses the fact that $\mathbf{WF}$ is reducible to $\mathcal{R}_{\mathrm{AP}}$. ∎
@@ 4. Complexity and the approximation property
This section is solely concerned with determining the complexity of two classes of Banach spaces having the AP. The proof of the result below is an adaptation of an argument from [17]. As above, we frame the result in terms of the Polish space $\mathcal{B}$ instead of the Effros-Borel space $\mathcal{SB}$.
@@
Hence we can define a bounded $\mathbb{Q}$-linear operator ${T}:{V}\to{\operatorname{span}(e_{i})_{i<R}}$ by setting $Te_{i}=w_{i}$, $i\in\omega$, and then extending to finite $\mathbb{Q}$-linear combinations of the $e_{i}$. By uniform continuity we extend it again to a bounded linear operator on $X_{\mu}$ satisfying $TX_{\mu}\subseteq\operatorname{span}(e_{i})_{i<R}$ and $\mu(T)\leqslant L$.